OpenStudy (anonymous):

Given: ∆BCA is a right triangle. Prove: a2 + b2 = c2 Which is NOT a justification for the proof? Substitution Addition Property of Equality Transitive Property of Equality Distributive Property of Equality

OpenStudy (anonymous):

Statement Justification Draw an altitude from point C to Line Segment AB Let Line Segment BC= a Line Segment CA= b Line Segment AB= c Line Segment CD= h Line Segment DB= x Line Segment AD= y y + x = c c over a is equal to a over y; c over b is equal to b over x a2 = cy; b2 = cx a2 + b2 = cy + b2 a2 + b2 = cy + cx a2 + b2 = c(y + x) a2 + b2 = c(c) a2 + b2 = c2

OpenStudy (anonymous):

^it wasnt supposed to say Justification, cuz there aint any... sry

OpenStudy (anonymous):

help?? please...

OpenStudy (anonymous):

I beleive its Transitive Property of Equality